My NolH8 AkYour Slote the deyrees uf freedum eor in ach of the folkowirng tests (h A researcher tasts haw narvaus pubic яр Akars (Mt in tmnt ถ『a emall, med im, ar lance aildianca. Twelve particinantsara randomly Aesignad to each ฏroup. A high school ccunselor has 11 tudents in cach of four clas rata how much thay lika thair teacher MinSitional Milli İlk eBcok -3 polnts Pristera Statc3 12.E.015 My Nones Ask Your
Answer #1

a)

degree of freedom =N-k =(16+16+16-3)=45

b)

degree of freedom =N-k =(12+12+12-3)=33

c)

degree of freedom =N-k =(11+11+11+11-4)=40